1)The milk and water in two vessels A and B are in the ratio 4:3 and 2:3 respectively. In what ratio,the liquids in both the vessels be mixed to obtain a new mixture in vessel C containing half milk and half water?
2)Two numbers A and B are such that the sum of 5% of A and 4% of B is 2/3rd of the sum of 6% of A and 8% of B. The ratio A:B is ?

Answer by mananth(16946)

Milk in 1 litre mixture of A = 4/7 litre.
Milk in 1 litre mixture of B = 2/5 litre.
Milk in 1 litre mixture of C = 1/2 litre.
By rule of alligation we have required ratio X:Y
X : Y
4/7 2/5

(Mean ratio)
(1/2)
(1/2 – 2/5) : (4/7 – 1/2)
1/10 1/1 4




Required ratio = X : Y = 1/10 : 1/14 = 7:5


Two numbers A and B are such that the sum of 5% of A and 4% of B is 2/3rd of the sum of 6% of A and 8% of B. The ratio A:B is ?
0.05A+0.04B= 2/3( 0.06A+0.08B)
0.15A +0.12B= 0.12A+0.16B
0.03A=0.04B
A/B = 4/3